# UrbanCAD: Towards Highly Controllable and Photorealistic 3D Vehicles for Urban Scene Simulation

### v1

@@ Abstract
Photorealistic 3D vehicle models with high controllability are essential for autonomous driving simulation and data augmentation. While handcrafted CAD models provide flexible controllability, free CAD libraries often lack the high-quality materials necessary for photorealistic rendering. Conversely, reconstructed 3D models offer high-fidelity rendering but lack controllability. In this work, we introduce UrbanCAD, a framework that pushes the frontier of the photorealism-controllability trade-off by generating highly controllable and photorealistic 3D vehicle digital twins from a single urban image and a large collection of free 3D CAD models and handcrafted materials. These digital twins enable realistic 360 rendering, vehicle insertion, material transfer, relighting, and component manipulation such as opening doors and rolling down windows, supporting the construction of long-tail scenarios. To achieve this, we propose a novel pipeline that operates in a retrieval-optimization manner, adapting to observational data while preserving flexible controllability and fine-grained handcrafted details. Furthermore, given multi-view background perspective and fisheye images, we approximate environment lighting using fisheye images and reconstruct the background with 3DGS, enabling the photorealistic insertion of optimized CAD models into rendered novel view backgrounds. Experimental results demonstrate that UrbanCAD outperforms baselines based on reconstruction and retrieval in terms of photorealism. Additionally, we show that various perception models maintain their accuracy when evaluated on UrbanCAD with in-distribution configurations but degrade when applied to realistic out-of-distribution data generated by our method. This suggests that UrbanCAD is a significant advancement in creating photorealistic, safety-critical driving scenarios for downstream applications.
@@ 1 Introduction
Photorealistic driving simulators have gained great attention for providing a safe and cost-effective way to evaluate driving algorithms  [68, 69, 62]. Digital twins of vehicles, representing key traffic participants, are essential for these simulators. Since simulators must assess driving algorithms in both common and rare, long-tailed scenarios, the vehicles within these simulators must exhibit both photorealism and controllability. This enables the manipulation of features such as pose, materials, door status, and wheel rotation, facilitating the creation of realistic yet uncommon scenarios for comprehensive evaluation.
@@
Classical game-engine-based driving simulators using handcrafted CAD models to represent vehicles like CARLA [16] achieve high controllability, yet exhibit a noticeable domain gap with the real world. Considering the easy-to-access real-world urban images, photorealistic simulation from these images provides an alternative solution to bridge this gap and scales to diverse scenarios. This direction has gained significant attention with the rapid development of the neural rendering techniques [22, 45, 52, 57, 65, 79]. While these methods eliminate domain gap and enable photorealistic rendering, they fall short in the controllability of the reconstructed vehicles. In particular, the vehicles are usually observed from limited viewpoints in urban scenes, impeding the photorealistic rendering of occluded regions. Although reconstructing the full vehicle from partial observations is possible by leveraging prior knowledge [47, 43], the quality of the reconstructed unobserved regions remains unsatisfactory. Furthermore, existing reconstruction-based methods rarely support control over material, lighting, and individual vehicle components. [62] employs CAD models of cars as shape priors and reconstructs vehicles with controllable wheels, appearance, and scene lighting using differentiable rendering. However, this approach still provides limited control over other vehicle components and yields unsatisfactory geometry results. Motivated by these observations, we seek to push the frontier of the photorealism-controllability trade-off.
@@
We move towards this objective by introducing a novel framework that automatically produces 3D vehicles with photorealistic appearance, fine-grained geometric details, and high controllability, including part-level control, from a single urban image and a collection of free 3D CAD models and handcrafted materials. This framework operates within a retrieval-optimization paradigm, performing both CAD retrieval and retrieval-based material optimization. Specifically, we represent vehicles using handcrafted CAD models, which offer high controllability due to disentangled designs, particularly part-disentangled geometry for component editing. This disentangled geometry results from the manual construction process, where engineers typically build vehicle CAD models component by component and integrate them into complete vehicles. Existing reconstruction-based methods [62, 26, 43] that use CAD models for training focus only on their appearance or overall geometry, neglecting their disentangled geometry design, which results in a loss of part controllability. In contrast, our image-based CAD retrieval is employed to preserve flexible controllability and fine-grained geometry by maintaining the representation of original handcrafted CAD models. We then retrieve part-aware material priors via foundation models and perform part-aware material optimization using physics-based differentiable rendering to enhance photorealism. In addition, given multi-view fisheye and perspective images of the background scene, we propose a fisheye-based spatially varying lighting estimation method to realistically render the optimized CAD model and reconstruct the background using 3D Gaussian Splatting [81] for high-fidelity novel view background synthesis. The integration of these renderings results in photorealistic novel view synthesis and versatile controllability over foreground vehicles. Note that our CAD retrieval and optimization module can process single-view images, allowing our method to work with diverse data sources, including internet images and those from dynamic urban scenes, which may lack high-quality multi-view observations. Additionally, recovering full vehicle materials from single-view images is feasible, as vehicles typically have symmetric material assignments, which our method exploits.
@@
Note that CAD retrieval and material optimization have been explored in several studies [17, 59, 31]. The most relevant work to ours is PhotoScene [71], which assumes that each object possesses a single material like furniture. In contrast, our pipeline accommodates objects with multiple materials, such as vehicles, through part-aware material retrieval and optimization. Additionally, while PhotoScene relies on manually retrieved CAD models or low-quality geometry from 3D reconstruction, our approach minimizes manual intervention and provides flexible control with fine-grained geometry through CAD retrieval. Further, we combine the strengths of CAD and implicit 3D representations, supporting photorealistic novel view synthesis and flexible editing. Finally, we validate our value through downstream applications for autonomous driving.
@@ 3.3 Material Optimization
Part-Aware Material Optimization. We also segment the components in the reference view using Grounded SAM and optimize the corresponding material of the CAD model to align with the reference view. Since there is no exact correspondence between rendered and reference pixels, we use a part-level loss $\mathbf{\ell}_{stat}$ by minimizing the difference between the mean and variance of the corresponding parts following  [71]. To match the patterns of the reference view, we use a masked VGG loss $\ell_{vgg}$ using Gram matrices [20] to enhance visual similarity. To further align the color of the reference vehicles, we incorporate a masked RGB loss $\ell_{rgb}$ on the overlap region between components in the reference view and CAD model rendering. Please see the supplementary Section 7.5 for details. The total loss optimizes the parameters of the material graphs with backpropagation. Note that optimizing special materials with ambiguous colors, such as glass, through differentiable rendering is not ideal. Instead, directly assigning base materials via our part-level material retrieval can produce satisfactory results, as demonstrated in our experiments.
@@ 4.2 Background Reconstruction and Compostion
Since autonomous driving simulators require free navigation within the scene, we integrate our method with novel view synthesis (NVS) to enable this functionality. Given input background videos, we use the 3D Gaussian Splatting method  [81] to reconstruct the environment and render background images from novel views. Subsequently, we render the foreground vehicle using Blender [2] and blend it with the background images via alpha composition. Specifically, we position the vehicle models generated in Section 3 in Blender according to a target 3D position. A virtual plane is utilized to account for shadow effects based on the estimated ground plane. Using the previously estimated HDR environment map, we render the vehicles and composite the rendered vehicles with the background images. Please refer to the supplementary Section 7.7 for details.
@@ 5.1 Experiment Setup
Baselines. We compare our approach with various types of methods: (1) Single-view reconstruction method using the conditional implicit function: PixelNeRF [73]. (2) Single-view generation method using diffusion prior: Wonder3D [43]. (3) Single-view reconstruction method using large reconstruction model: LRM [26]. (4) Multi-view reconstruction method using 3DGS: HUGS [81]. (5) Mesh texturing methods using the generative model: LatentPaint [46] and Paint3D [75]. (6) Mesh texturing method using optimizable procedural graph: PhotoScene [71]. Note that we use our CAD retrieval module (Section 3.1) to retrieve CAD models before mesh texturing with LatentPaint [46], Paint3D [75] and PhotoScene [71] (marked as “LatentPaint”, “Paint3D” and “PhotoScene”). We also investigate the performance when directly using our CAD retrieval module without material retrieval and optimization (marked as UrbanCAD (w/o opt.) ).
@@ 7.2 Pose Matching
Figure 11: Quality results of part-recognition based on randomly colored material design (top) and retrieved CAD renderings without ControlNet augmentation (bottom) with the text prompt of ”windows”.
@@ 7.5 Material Optimization
In our experiment, we set the $\lambda_{stat}$ to 0.1, the $\lambda_{vgg}$ to 1, the $\lambda_{rgb}$ to 1. Note that spatially varying roughness parameters are difficult to optimize from single-view images due to limited highlight observations. Handcrafted procedural material graphs provide photorealistic spatially varying effects, so the roughness parameters of the retrieved material prior are fixed during optimization, as in [71]. Besides, we observe two types of materials with distinct spatially varying effects in car bodies depending on whether the vehicles are painted or not, as shown in Fig. 12. To best fit the observation, we select the corresponding car body material prior via the user interface.
@@ 9.1 Component Editing
Our produced 3D vehicle models support easy component editing mainly due to the handcrafted disentangled geometry as shown in Fig. 10. Note that complete component editing requires human effort for animation, such as setting joint types and parameters in Blender. Additionally, some retrieved handcrafted CAD models may have merged geometry, for example, the four wheels are merged in one mesh. For these cases, simply hiding other vehicle components and entering the edit mode to separate the wheels by selection in the Blender can solve the problem with small manual efforts. However, we notice that some vehicle CAD models have been post-processed by geometry merging, such as the vehicles in ShapeNet [10], which means the loss of part controllability. Fortunately, most handcrafted vehicle CAD models in the Objaverse still preserve part controllability without being post-processed, and many post-processed CAD models still have disconnected geometry, which can be manually separated by Blender “Separate Selection” operation after selecting connected geometry (“Select Linked” function in “Select” menu). Besides, more corner case results are displayed in Figure 15 thanks to the representation of CAD models. In addition to the editing results mentioned earlier, we can generate more scenes, using the powerful physical simulation effects in Blender. By assigning physics properties to the vehicle model, we can create collision scenes or even simulate car accidents in Blender.

### v2

@@ Abstract
Photorealistic 3D vehicle models with high controllability are essential for autonomous driving simulation and data augmentation. While handcrafted CAD models provide flexible controllability, free CAD libraries often lack the high-quality materials necessary for photorealistic rendering. Conversely, reconstructed 3D models offer high-fidelity rendering but lack controllability. In this work, we introduce UrbanCAD, a framework that generates highly controllable and photorealistic 3D vehicle digital twins from a single urban image, leveraging a large collection of free 3D CAD models and handcrafted materials. To achieve this, we propose a novel pipeline that follows a retrieval-optimization manner, adapting to observational data while preserving fine-grained expert-designed priors for both geometry and material. This enables vehicles’ realistic 360 rendering, background insertion, material transfer, relighting, and component manipulation. Furthermore, given multi-view background perspective and fisheye images, we approximate environment lighting using fisheye images and reconstruct the background with 3DGS, enabling the photorealistic insertion of optimized CAD models into rendered novel view backgrounds. Experimental results demonstrate that UrbanCAD outperforms baselines in terms of photorealism. Additionally, we show that various perception models maintain their accuracy when evaluated on UrbanCAD with in-distribution configurations but degrade when applied to realistic out-of-distribution data generated by our method. This suggests that UrbanCAD is a significant advancement in creating photorealistic, safety-critical driving scenarios for downstream applications.
@@ 1 Introduction
Photorealistic driving simulators have gained great attention for providing a safe and cost-effective way to evaluate driving algorithms  [68, 69, 62]. Digital twins of vehicles, representing key traffic participants, are essential for these simulators. Since simulators must assess driving algorithms in both common and rare, long-tailed scenarios, the vehicles within these simulators must exhibit both photorealism and controllability.
@@
Classical driving simulators based on game engines, such as CARLA [16], use handcrafted CAD models to represent vehicles. While offering high controllability, they suffer from a significant domain gap compared to the real world. Leveraging easily accessible real-world urban images, photorealistic simulation offers a scalable solution to bridge this gap across diverse scenarios. This direction has gained wide attention with advances in neural rendering techniques [44, 51, 57, 65, 56]. While these methods close the domain gap and enable photorealistic rendering, they lack fine-grained control over the reconstructed vehicles. [62] employs CAD models of cars as shape priors and reconstructs vehicles with controllable wheels, appearance, and scene lighting using differentiable rendering. However, it still provides limited control over other vehicle components and yields suboptimal geometry. In addition, vehicles are usually observed from limited viewpoints in urban scenes, impeding the photorealistic rendering of occluded regions. Although prior knowledge can help reconstruct unobserved parts [46, 42], the quality of these regions remains unsatisfactory. Furthermore, complex material physical properties such as opacity and metallic reflectance, pose significant challenges for differentiable rendering, especially in our single-view observation setting. In contrast, handcrafted material libraries like Adobe Material Library [1] provide materials with complex physical properties. Motivated by these observations, we seek to push the frontier of the photorealism-controllability trade-off.
@@
We move towards this objective by introducing a novel framework that automatically produces 3D vehicles with photorealistic appearance, fine-grained geometric details, and high controllability, including part-level control, from a single urban image and a collection of free 3D CAD models and handcrafted materials. This framework operates within a retrieval-optimization paradigm, performing both CAD retrieval and retrieval-based material optimization. The key idea is to retain the existing detailed expert-designed priors including part-level controllability and complex material physical properties, while refining the appearance to fit the observational data. Specifically, we first retrieve vehicles’ geometries represented by handcrafted CAD models, which offer high controllability due to disentangled designs, particularly part-disentangled geometry for component editing. Existing reconstruction-based methods [62, 26, 42] that use CAD models for training focus only on their appearance or overall geometry, neglecting their disentangled geometry design, which results in a loss of part controllability. Then, we retrieve vehicles’ materials represented by optimizable, manually designed procedural graphs, which offer photorealistic material properties—such as opacity and roughness—that are difficult to optimize directly. However, for vehicles that have multiple types of materials, accurately retrieving and assigning the part-aware material priors are not trivial. Previous works [71, 55] typically retrieve material priors based on visual similarity, which is inaccurate for materials with complex physical properties, and primarily focus on objects with a single material. To address this, we retrieve part-aware materials based on semantic meanings via foundation models and assign materials through a ControlNet-based recognition method informed by the retrieved material designs. Finally, we perform part-aware material optimization using physics-based differentiable rendering to align the appearance with the input image. In addition, given multi-view fisheye and perspective images of the background scene, we propose a fisheye-based spatially varying lighting estimation method to realistically render the optimized CAD model and reconstruct the background using 3D Gaussian Splatting [82] for high-fidelity novel view background synthesis. The integration of these renderings results in photorealistic novel view synthesis and versatile controllability over foreground vehicles.
@@ 3.3 Material Optimization
Part-Aware Material Optimization. We also segment the components in the reference view using Grounded SAM and optimize the corresponding material of the CAD model to align with the reference view. Since there is no exact correspondence between rendered and reference pixels, we use a part-level loss $\mathbf{\ell}_{stat}$ by minimizing the difference between the mean and variance of the corresponding parts following  [71]. To match the patterns of the reference view, we use a masked VGG loss $\ell_{vgg}$ using Gram matrices [20] to enhance visual similarity. To further align the color of the reference vehicles, we incorporate a masked RGB loss $\ell_{rgb}$ on the overlap region between components in the reference view and CAD model rendering. Please see the supplementary Section 7.5 for details. The total loss optimizes the parameters of the material graphs with backpropagation. Note that optimizing complex material physical properties from single view images, such as opacity and roughness through differentiable rendering is not ideal. Instead, we only optimize the albedo of the retrieved metal and rubber materials to align with the input image and directly assign the retrieved glass material to the car window without further optimization, which can produce satisfactory results, as demonstrated in our experiments.
@@ 4.1 Environment Lighting Estimation
Ref.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Rot.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ref.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Rot.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PixNeRF [73] Wonder3D [42] LRM [26] HUGS [82] Ours (w/o opt.) LatentPaint [45] Paint3D [75] PhotoScene[71] Ours GT
@@ 4.2 Background Reconstruction and Compostion
Since autonomous driving simulators require free navigation within the scene, we integrate our method with novel view synthesis (NVS) to enable this functionality. Given input background videos, we use the 3D Gaussian Splatting method  [82] to reconstruct the environment and render background images from novel views. Subsequently, we render the foreground vehicle using Blender [2] and blend it with the background images via alpha composition. Specifically, we position the vehicle models generated in Section 3 in Blender according to a target 3D position obtained from annotations or trajectory generation methods [77]. A virtual plane is utilized to account for shadow effects based on the estimated ground plane. Using the previously estimated HDR environment map, we render the vehicles and composite the rendered vehicles with the background images. Please refer to the supplementary Section 7.7 for details.
@@ 5.1 Experiment Setup
Baselines. We compare our approach with various types of methods: (1) Single-view reconstruction method using the conditional implicit function: PixelNeRF [73]. (2) Single-view generation method using diffusion prior: Wonder3D [42]. (3) Single-view reconstruction method using large reconstruction model: LRM [26]. (4) Multi-view reconstruction method using 3DGS: HUGS [82]. (5) Mesh texturing methods using the generative model: LatentPaint [45] and Paint3D [75]. (6) Mesh texturing method using optimizable procedural graph: PhotoScene [71]. Note that we use our CAD retrieval module (Section 3.1) to retrieve CAD models before mesh texturing with LatentPaint [45], Paint3D [75] and PhotoScene [71] (marked as “LatentPaint", “Paint3D" and “PhotoScene"). We also investigate the performance when directly using our CAD retrieval module without material retrieval and optimization (marked as UrbanCAD (w/o opt.) ).
@@ 5.2 Photorealism Quality
Ref. Ours. Ref. Ours. Ref. Ours. Ref. Ours.
Figure 6: Recovery from partial observation.
Robustness to occlusion. We present more results based on partial observations in Fig. 6. We find SAM is robust to occlusion, likely due to its training on occluded images. Our CAD retrieval module demonstrates resilience to occlusion as well, as it utilizes a 3D encoder aligned with CLIP’s latent space to encode CAD models and retrieve them based on semantic instead of visual similarity. In addition, our material optimization module produces reliable results by leveraging part-aware regularization between corresponding semantic regions, eliminating the need for precise pixel-level alignment.
@@ 7.2 Pose Matching
Figure 12: Quality results of part-recognition based on randomly colored material design (top) and retrieved CAD renderings without ControlNet augmentation (bottom) with the text prompt of "windows".
@@ 7.5 Material Optimization
In our experiment, we set the $\lambda_{stat}$ to 0.1, the $\lambda_{vgg}$ to 1, the $\lambda_{rgb}$ to 1. Note that spatially varying roughness parameters are difficult to optimize from single-view images due to limited highlight observations. Handcrafted procedural material graphs provide photorealistic spatially varying effects, so the roughness parameters of the retrieved material prior are fixed during optimization, as in [71]. Besides, we observe two types of materials with distinct spatially varying effects in car bodies depending on whether the vehicles are painted or not, as shown in Fig. 13. To best fit the observation, we recommend selecting the corresponding car body material prior via the user interface.
@@ 9.1 Component Editing
Our produced 3D vehicle models support easy component editing mainly due to the handcrafted disentangled geometry as shown in Fig. 11. Note that complete component editing requires human effort for animation, such as setting joint types and parameters in Blender. Additionally, some retrieved handcrafted CAD models may have merged geometry, for example, the four wheels are merged in one mesh. For these cases, simply hiding other vehicle components and entering the edit mode to separate the wheels by selection in the Blender can solve the problem with small manual efforts. However, we notice that some vehicle CAD models have been post-processed by geometry merging, which means the loss of part controllability. Fortunately, most handcrafted vehicle CAD models in the Objaverse still preserve part controllability without being post-processed, and many post-processed CAD models still have disconnected geometry, which can be manually separated by Blender “Separate Selection" operation after selecting connected geometry (“Select Linked" function in “Select" menu). Besides, more corner case results are displayed in Figure 16 thanks to the representation of CAD models. In addition to the editing results mentioned earlier, we can generate more scenes, using the powerful physical simulation effects in Blender. By assigning physics properties to the vehicle model, we can create collision scenes or even simulate car accidents in Blender.
@@ 10.3 Failure Cases
[…]
Ours 0.052 0.636
Wonder3D 0.058 0.588
Table 6: Geometry quality
10.4 Geometry quality.
We randomly select 30 vehicles from the ShapeNet dataset, encompassing various types, and retrieve their corresponding models from the Objaverse dataset. We report the Chamfer Distance and Volume IOU in Table 6. Our retrieved models’ geometry quality surpasses the reconstruction baseline, Wonder3D [42], as our retrieved CAD models often exhibit better geometry quality in unobservable regions.